OpenStudy (anonymous):

How much water must be added to 80 liters of 60% alcohol solution in order to get 35% alcohol solution ?

OpenStudy (anonymous):

Calculate the number of litres of alcohol first. 60% of 80 litres= 0.6*80= 48 liters. To get a 35% alcohol solution, that means 48 liters is 35% of the total volume of the solution. Since 48 is the 35% find 100%, 48/35 * 100= 137liters. How much water must be added? 137 - 80 = 57 liters.